### v2

# Computing Anharmonic Infrared Spectra of Polycyclic Aromatic Hydrocarbons Using Machine-Learning Molecular Dynamics

Abstract
Polycyclic aromatic hydrocarbons (PAHs) are key contributors to interstellar aromatic infrared (IR) bands. However, current spectral databases for IR emission analysis are limited by the omission of vibrational anharmonicity and temperature effects, primarily because of the high computational cost of conventional quantum chemical calculations (QCCs). In this work, we present a machine learning-based molecular dynamics (MLMD) approach that efficiently computes anharmonic IR spectra while incorporating temperature effects. MLMD achieves predictive accuracy comparable to that of QCCs but with significantly reduced computational cost, scaling linearly with the number of atoms in the system. We applied MLMD to calculate the anharmonic spectra of 1 704 PAHs in the NASA Ames PAH IR Spectroscopic Database with up to 216 carbon atoms, demonstrating its capability for high-throughput spectral calculations of large molecular systems. Our results highlight MLMD’s potential to enable the development of extensive molecular spectral datasets, enhancing data-driven analyses of astronomical IR spectra, particularly in anticipation of upcoming data from the James Webb Space Telescope.
astronomical data bases: miscellaneous; software: data analysis; ISM: molecules; infrared: ISM
I Introduction
In the mid-1970s, distinct infrared (IR) signals were identified in the universe, notably at wavelengths of 3.3, 6.2, 7.7, 8.6, 11.2, and 12.7 $\upmu$m. These signals were later linked to the vibrations of C-H and C-C bonds in aromatic molecules, known as aromatic IR bands (AIB) (Leger & Puget, 1984; Allamandola et al., 1985). Such emissions are primarily attributed to polycyclic aromatic hydrocarbons (PAHs) after the absorption of ultraviolet (UV) photons. PAHs are recognized as key contributors to the evolution of the interstellar medium, influencing processes such as gas heating, ionization balance, and star formation (Tielens, 2008). Determining the population, composition, size distribution, charge state, and chemical structures of interstellar PAHs provides valuable insight into the physical conditions of their host galaxies. However, this remains a significant challenge because of the structural diversity of PAHs. Variations in carbon-ring arrangements, side groups, substitutions, charge states, and isotopologues lead to distinct IR emission characteristics, each governed by specific vibration modes (Peeters et al., 2021).
In order to tackle these challenges, quantum chemical calculations (QCCs) are now an indispensable aid in AIB analysis. Using QCCs, publicly available spectral databases have been established for AIB analysis, such as the NASA Ames PAH IR Spectroscopic Database (PAHdb) (Mattioda et al., 2020; Bauschlicher et al., 2018; Boersma et al., 2014). The availability of these databases enables extensive data-driven analysis of AIBs, providing a more robust interpretation of their features (Boersma et al., 2013, 2015; Maragkoudakis et al., 2020; Sadjadi et al., 2015; Cami, 2011; Shannon & Boersma, 2019; Ricca et al., 2021). Furthermore, these databases have facilitated the analysis of astrochemical components using machine learning (ML) models, demonstrating significant computational efficiency (Kovács et al., 2020) and helping to uncover relationships between molecular structures and spectra (Meng et al., 2021, 2023). Despite these advancements, current theoretical spectral databases face three primary limitations:
  • Harmonic oscillator approximations: Most spectra neglect anharmonic vibrations, leading to significant discrepancies with experimental observations (Mackie et al., 2015, 2018a; Maltseva et al., 2016, 2018; Lemmens et al., 2019).
  • Neglect of temperature effect: AIB features are typically attributed to IR emission from far-UV-pumped PAHs (Leger & Puget, 1984), where an absorbed UV photon can elevate the effective molecular temperature to several hundred or even thousands of kelvin, an effect that is not taken into account in most theoretical spectral data.
  • Predominantly featuring small molecules: For example, a significant portion of PAHs in PAHdb consists of fewer than 35 carbon atoms, while larger PAHs ($\sim$50 carbon atoms) are more likely sources of AIBs (Sellgren, 1984; Allamandola et al., 1989; Bauschlicher et al., 2009; Maragkoudakis et al., 2020)
These challenges arise mainly from the high computational cost of QCCs. In particular, the computation time for anharmonic molecular spectra using second-order vibrational perturbation theory (VPT2) is prohibitively long, typically only accommodating molecules with $N_{\text{C}}<25$, where $N_{\text{C}}$ represents the number of carbon atoms in the system (Mackie et al., 2016; Esposito et al., 2024a). The cost of incorporating temperature effects into these calculations using the Wang-Landau random walk algorithm is even more prohibitive, being orders of magnitude higher due to the need for extensive sampling (Chen, 2018; Chen et al., 2019). Facing the vast diversity of PAH species, these limitations significantly hinder the accurate interpretation of observational data, posing substantial challenges in the era of advanced observational facilities such as the James Webb Space Telescope (JWST) (Boersma et al., 2023; Ricca et al., 2024).
As a result, there is an urgent need for a cost-effective and accurate method that incorporates both anharmonic and temperature effects in the construction of spectral databases to improve our understanding of the origin of AIBs. In this work, we propose a ML-based molecular dynamics (MLMD) approach for computing PAH IR spectra. This method replaces computationally expensive electronic structure calculations with far more efficient ML-based ones. More importantly, it explicitly accounts for anharmonic vibrations and temperature effects. We show that the MLMD approach can reproduce experimental spectra with accuracy comparable to state-of-the-art VPT2 QCCs, while significantly reducing computational time, achieving a scaling approximately linear in $N_{\text{C}}$.
II Methods
Molecular dynamics (MD) is a powerful computational method that simulates the time-dependent behavior of atomic and molecular systems at a specified temperature by integrating Newton’s equations of motion. This technique uses a potential energy surface (PES) to characterize interatomic interactions and energy changes within the system. When calculating the vibrational spectra of molecules, MD presents significant advantages by explicitly accounting for anharmonic effects, such as band combination, overtones, and mode coupling. This comes at the cost of not capturing some quantum effects whenever normal modes are likely to be found near their ground state. Beyond this fundamental consideration, classical MD has historically faced technical challenges related to accuracy and transferability. The empirical force fields used to construct the PES are often parameterized for specific molecules, which may hinder their performance when applied to others. Furthermore, although polarizable models exist (Bedrov et al., 2019), classical MD generally does not account for the distribution of electrons and, therefore, cannot inherently deal with dipole moments. To address these challenges, ab initio molecular dynamics (AIMD) can be used, where the motion of nuclei is described classically but the force contributions from the electrons are calculated quantum mechanically. However, like the QCC-based VPT2 method, AIMD is computationally expensive, imposing significant constraints on the maximum size of the systems studied.
A solution is to substitute the majority of electronic structure calculations in AIMD with more cost-effective ML computations. Using data points derived from QCCs, ML models can be trained to construct the PES and predict the charge distribution across various molecular configurations. This MLMD methodology has been proven to be highly efficient in previous studies focusing on vibration of diverse molecular types (Gastegger et al., 2017; Zhou et al., 2021; Du et al., 2024; Xu et al., 2024; Schmiedmayer & Kresse, 2024).
Figure 1: Schematic representation of the workflow for computing anharmonic IR spectra.
In this study, we build two distinct ML models: a neural-network force field (NNFF) for the construction of PES, and an electron-passing neural network (EPNN) model to predict the molecular dipole moment $\mathbf{p}$. As depicted in Figure 1, our workflow for computing the IR spectra consists of two main phases. The first is a training phase (blue arrows) that includes:
  1. 1.
    Classical MD simulations of PAHs to generate conformations;
  2. 2.
    QCCs to obtain their energies, forces, and dipole moments;
  3. 3.
    Training of the NNFF and EPNN models using these data.
The second phase is prediction (orange arrows), including:
  1. 1.
    MLMD to obtain conformations of vibrating PAHs via NNFF;
  2. 2.
    Prediction of $\mathbf{p}$ for every conformation using EPNN;
  3. 3.
    Spectrum computation by evaluating the time evolution of $\mathbf{p}$.
The predicted $\mathbf{p}$ is not used as an input for the calculation of energies and forces and therefore does not determine the trajectories, since the interatomic interactions determined by the charge density distribution are already accounted for by the NNFF. A detailed description of the components is provided in the next subsection.
II.1 Training Data
The chemical structures of 687 neutral PAH molecules were extracted from the theoretical dataset of PAHdb (v3.2, Mattioda et al. (2020)) to construct the training datasets. The selected PAHs encompass a diverse array of chemical structures that are of astronomical significance. Classical MD simulations were performed using a custom code to generate random configurations of the atoms in each molecule at 300 K, using the adaptive interatomic reactive empirical bond order potential (Stuart et al., 2000). In these simulations, the molecules reached thermal equilibrium in 200 ps using a canonical Nose-Hoover thermostat, with a time step of 0.5 fs. Ten extended structures were extracted from the atomic trajectories of each molecule. As these geometries do not necessarily fall near equilibrium configurations, random Gaussian perturbations were applied to the ground-state atomic coordinates to create 15 more configurations for each molecule to improve the generalizability of the model. Finally, in total, 17 175 configurations (also called molecular conformations) were generated.
Subsequently, single-point electronic structure QCCs were performed for each of these 17 175 configurations to calculate the potential energy, the forces in the nuclei, and the dipole $\mathbf{p}$, within the framework of density functional theory (DFT) at the B3LYP/4-31G level, as implemented in the Gaussian 16 software package (Frisch et al., 2016). This level of theory offers a favorable balance between accuracy and computational efficiency for large PAHs (Stephens et al., 1994), given its extensive application in PAH IR spectrum studies (Bauschlicher & Bakes, 2000; Ricca et al., 2012). We note that the basis set does not include polarization functions, which may introduce more uncertainties than larger basis functions (e.g 6-31g*) for PAHs containing nitrogen (Ricca et al., 2021, 2024). However, this choice was made to ensure computational efficiency, which is crucial for large-scale sampling. Through these calculations, we generated two datasets for training the NNFF and EPNN models, respectively. The first dataset comprises 17 175 energy and 2 274 825 force data points, while the second dataset includes 17 175 molecular dipole moments.
II.2 NNFF
The NNFF used in this study is based on NeuralIL (Montes-Campos et al., 2021; Carrete et al., 2023), a refinement of the descriptor-based template introduced by Behler (Behler & Parrinello, 2007). NeuralIL employs spherical Bessel descriptors (Kocer et al., 2020) to encode the atomic configurations in a translationally and rotationally invariant fashion, as illustrated in Figure 2a. We set the cutoff radius of the descriptors to 3.8 Å, with a maximum radial order of 6, ensuring effective capture of local geometric information and sufficient resolution to describe atomic interactions within the molecule.
Figure 2: (a) Schematic of the operating principles of each kind of neural network (NNFF and EPNN) during inference. (b) NNFF-predicted atomic forces and (c) EPNN-predicted dipole moments vs. the DFT-calculated reference values for 3 580 tested PAH configurations.
NNFF utilizes a deep residual network architecture (ResNet) (He et al., 2016a), implemented within the JAX framework (Bradbury et al., 2018). A ResNet employs skip connections to learn the residual function between the input and output layers, effectively addressing the data degradation problem that arises with increased network depth (He et al., 2016b). This feature is particularly beneficial for modeling complex molecules such as large PAHs, as it enables the capture of deep dependencies and interaction patterns between atoms through multilevel learning, which ultimately leads to more accurate predictions of physicochemical properties (Zhang et al., 2024; Xue et al., 2024).
In this study, the NNFF is trained based on the previously mentioned DFT-calculated dataset of energies $E$ and atomic forces $\mathbf{f}$, with the loss function $L$ defined as a weighted sum of the differences between the predicted and reference values:
$\begin{split}L&=0.99\sum^{N_{\text{mol}}}\left\langle\frac{0.2}{N_{\text{atom}}}\sum^{N_{\text{atom}}}\mathrm{ln}\left[\mathrm{cosh}\left(\frac{\left\|\mathbf{f}_{\text{pred}}-\mathbf{f}_{\text{ref}}\right\|_{2}}{0.2}\right)\right]\right\rangle\\ &+0.01\sum^{N_{\text{mol}}}\left\langle 0.02\ \mathrm{ln}\left[\mathrm{cosh}\left(\frac{E_{\text{pred}}-E_{\text{ref}}}{N_{\text{atom}}\times 0.02}\right)\right]\right\rangle,\end{split}$ (1)
where $N_{\text{mol}}$ denotes the total number of molecules, and ${N_{\text{atom}}}$ represents the total number of atoms in a molecule. The subscripts ${}_{\text{pred}}$ and ${}_{\text{ref}}$ denote the values predicted by NNFF and DFT, respectively. $\left\|\mathbf{f}_{\text{pred}}-\mathbf{f}_{\text{ref}}\right\|_{2}$ represents the Euclidean norm of the vector difference between the predicted force and the reference force. The model adopts a core-width sequence of 64:32:16:16 and employs the fully non-linear VeLO optimization technique (Metz et al., 2022) for 600 epochs of training to enhance convergence (Carrete et al., 2023).
To evaluate the performance of the model, we conducted tests on a set of configurations for 3 580 PAHs that were excluded from the training set. The results presented in Figure 2b demonstrate that the NNFF model accurately reproduces the forces calculated by DFT, achieving a coefficient of determination (${R^{2}}$) of approximately 0.99 and a mean root mean square error (RMSE) of 0.00117 hartree/bohr (see Table 1), indicating strong generalizability to unseen samples. This capability allows the NNFF model to effectively replace DFT electronic structure calculations in AIMD, providing robust support for subsequent IR spectrum calculations.
Table 1: Average RMSE of the predicted energy and force from the NNFF, and of the predicted molecular dipole moments from the EPNN, for samples in the training, validation and test datasets.
Train Validation Test
Energy (hartree per atom) 0.315 0.323 0.319
Force (hartree/bohr) 0.00118 0.00117 0.00115
Dipole (debye) 0.0550 0.0545 0.0556
II.3 EPNN
Predicting the molecular dipole moment $\mathbf{p}$ in a picture of discrete atoms presents a significant challenge due to its dependence on the distribution of atomic partial charges (PC), for which no unique determination method exists. Various PC partitioning schemes can produce different $\mathbf{p}$ for the same molecule. To avoid this issue, we directly train neural networks to predict $\mathbf{p}$, bypassing the need to explicitly calculate PCs as training data, following the approach of Gastegger et al. (2017). However, we utilize a different neural network architecture for enhanced predictive ability. Specifically, we adopt a message-passing neural network known as the EPNN (Metcalf et al., 2021) that guarantees the conservation of charge. In order to achieve charge conservation, the EPNN adopts a graph neural network architecture and ensures anti-symmetry of the updates with respect to permutation of the input indices. In the present implementation of the EPNN the node states are extended by concatenation with the spherical Bessel descriptor for each individual atom. Charge conservation is preserved as these descriptors are not modified in the update phase.
The EPNN was trained over 400 epochs using VeLO, employing the aforementioned DFT-calculated $\mathbf{p}$ dataset. The loss function used for training is defined as
$L=\sum\limits_{\alpha=1}^{3}\sum_{i=1}^{N_{\text{mol}}}\ln\left[\cosh\left(\frac{{{p}^{\left(\alpha\right)}_{i,\text{pred}}-{p}^{\left(\alpha\right)}_{i,\text{ref}}}}{N_{\text{atom}}}\right)\right],$ (2)
where $p^{\left(\alpha\right)}_{i,\text{pred}}$ and $p^{\left(\alpha\right)}_{i,\text{ref}}$ represent the predicted and reference values of the $\alpha$-th Cartesian component of $\mathbf{p}$ for the $i$-th molecule. To evaluate the performance of the model, we predicted the $\mathbf{p}_{\text{ref}}$ of 3 580 PAH conformations excluded from the training set. As shown in Figure 2c, there is strong agreement between the predicted and reference values, with a ${R^{2}}$ reaching 0.97 and a mean RMSE of 0.0545 debye, as shown in Table 1.
II.4 Anharmonic IR spectrum
To compute the anharmonic IR spectrum of a PAH, MD simulations were performed to simulate its finite-temperature dynamics based on atomic forces predicted by the trained NNFF. We used the MD implementation in the Atomic Simulation Environment (ASE) (Larsen et al., 2017). Initially, the molecule reaches thermal equilibrium in 200 ps in a canonical ensemble (NVT) at a target temperature controlled by the Nose-Hoover thermostat, with a timestep of 0.5 fs. The system was then further simulated in a microcanonical ensemble (NVE) for 200 ps, recording atomic configurations at every 1.0 fs during vibration.
These configurations serve as input to the EPNN, which is used to compute $\mathbf{p}(t)$ at each time step of the vibration. The dipole time autocorrelation function $\langle\mathbf{p}(0)\cdot\mathbf{p}(t)\rangle$, characterizing the time evolution of $\mathbf{p}$, is then subjected to a Fourier transform to obtain the IR intensity(berens1981):
[…]
where $\omega$ denotes the angular frequency, $k_{\text{B}}$ represents the Boltzmann constant, $\hbar$ is the reduced Planck constant, $c$ is the speed of light, $T$ indicates the temperature, and $t$ is time.
II.5 Reference methods for comparison
To evaluate the accuracy of our model, we compared it with two established QCC methods to calculate the IR spectra of PAHs. The first method, designated as DFT harmonic, is a widely adopted hybrid DFT approach for harmonic spectrum calculations. In this method, the vibrational frequencies (normal modes) are calculated from the second derivatives of the potential energy with respect to the positions of the nuclei at the stationary point of the system. IR intensities are determined using the double harmonic approximation, which involves computing the derivatives of $\mathbf{p}$ with respect to the normal modes (Langhoff, 1996). The calculations were carried out at the B3LYP/4-31G level of theory using Gaussian 16, ensuring consistency with the QCCs employed in our training dataset. It is important to note that the assumption of a harmonic potential frequently underestimates fundamental frequencies, often necessitating empirical scaling factors to align the computational results with the experimental data (Bauschlicher & Langhoff, 1997). Despite the transferability issues introduced by these empirical scaling factors, this method remains the most commonly utilized method for molecular IR spectrum calculations due to its relatively low computational cost. For the sake of a fair comparison, we do not apply any scaling factors to the DFT harmonic results, as these factors should ideally be derived from fitting to experimental spectra. The DFT harmonic method was used to compute the harmonic spectra for 49 PAH molecules, for which the experimentally measured spectra are available in the PAHdb.
The second method used for comparison is the VPT2 anharmonic spectrum calculations combined with B3LYP/6-311+g* DFT, referred to here as DFT anharmonic. This approach represents the state-of-the-art for computing the molecular anharmonic IR spectra, and has been implemented in Gaussian 16. It requires the computation of the second, third, and fourth derivatives of the potential, which results in significant computational demands (Nielsen, 1951; Mackie et al., 2015). Due to this cost, we applied VPT2 to calculate the anharmonic spectra of 26 PAHs, with the maximum $N_{C}$ being 32, from the 49 PAHs used for the DFT harmonic method.
The 49 PAHs for which the experimental spectra are available in PAHdb (v3.2), ranging from 10 to 50 carbon atoms, exhibit a variety of chemical structures, as detailed in Appendix I. The experimental IR spectra were obtained using matrix isolation techniques at low temperatures (Hudgins & Allamandola, 1995; Hudgins & Sandford, 1998a, b, c; Hudgins & Allamandola, 1999; Mattioda et al., 2003, 2005, 2014, 2017). To facilitate comparison with experimental data, we employed theoretical spectra computed using MLMD at an effective temperature of 50 K, which accounts for molecular vibrations under experimental conditions (Esposito et al., 2024b). In the comparison between computed and experimental spectra, discrete infrared spectral lines were Lorentzian broadened with a full width at half maximum of $18\text{cm}^{-1}$.
III Results and discussion
To assess the precision of our predictions relative to the experimental data, Figure 3 presents the computed spectra for five PAHs of increasing sizes, ordered from top to bottom. The black curves represent the experimental spectra, while the red, green, and blue curves show the spectra computed using MLMD, DFT harmonic, and DFT anharmonic methods, respectively. As seen in the figure, the anharmonic spectra (represented by the red and blue curves) exhibit better agreement with the experimental data than the harmonic spectra (green), both in terms of frequency and intensity.
Figure 3: Comparison of IR spectra for five PAHs computed using the MLMD at 50 K(left, red), DFT harmonic (middle, green), and DFT anharmonic (right, blue) methods, alongside the experimentally measured spectra (black). The intensities are normalized to the maximum values across the four spectra for ease of comparison. The RMSE is computed from the original intensity values, with units of $10^{5}\text{cm}^{2}/\text{mol}$.
[…]
Figure 4 presents the distribution of errors in predicting the IR spectra of the 49 PAHs relative to the experimental data, for each of the three computational methods. The results clearly demonstrate that the harmonic spectra exhibit significantly poorer agreement with the experimental data compared to the anharmonic spectra, consistent with the findings of previous studies (Mackie et al., 2015; Maltseva et al., 2016; Lemmens et al., 2019). The precision of the MLMD and DFT anharmonic methods is comparable, with mean RMSE values of 0.363 and 0.360 ×$\times 10^{5}\text{cm}^{2}/\text{mol}$, respectively. For PAHs with fewer than 16 carbon atoms, the DFT anharmonic method generally outperforms MLMD. However, as the size of the PAHs increases, MLMD tends to provide more accurate predictions than DFT anharmonic.
Figure 4: Distribution of RMSE values for spectra predicted by the three computational methods, compared to the experimental spectra of the 49 PAHs. $N_{\text{C}}$ represents the number of carbon atoms in each molecule. Note that the anharmonic DFT calculation is not performed for large PAHs (Section II.5).
Despite having comparable accuracy, the computational efficiency of MLMD is significantly superior to that of the traditional DFT anharmonic method. Figure 5 shows the average CPU time, $t_{\text{cpu}}$, required to compute an anharmonic IR spectrum as a function of the molecular size. It is evident that the $t_{\text{cpu}}$ for MLMD is initially approximately five times shorter than that for the DFT anharmonic for the smallest PAHs. As the molecular size increases, the difference in $t_{\text{cpu}}$ grows, eventually reaching nearly two orders of magnitude for the largest PAH in the set of 49 molecules.
Figure 5: Average CPU time (run time $\times$ number of cores, on a logarithmic scale) required to compute the anharmonic IR spectrum of a PAH as a function of the number of carbon atoms, using the MLMD (red dots) and DFT anharmonic (blue dots) methods. The DFT anharmonic computations were parallelized on a server with 40 cores (Intel Xeon E5-2680 CPU at 2.40 GHz), while MLMD jobs were run sequentially on a CPU of the same model.
The trend of increasing CPU time with molecular size for MLMD is remarkably different from that of the DFT anharmonic method. To provide a quantitative illustration, a simple power-law fit of type $y=ax^{b}$ was applied to the data points in Figure 5, as indicated by the dashed curves. These best-fitting curves show that the DFT computation of anharmonic spectra for large PAHs becomes extremely time-consuming, with a computational scaling of $\sim N_{\text{C}}^{4}$ for the DFT anharmonic method. In contrast, MLMD demonstrates a significantly more efficient scaling of approximately $N_{\text{C}}^{0.7}$. This results in a computational time difference of approximately three orders of magnitude for a PAH with $N_{\text{C}}=50$. This enhanced efficiency makes MLMD well suited for high-throughput computation of molecular anharmonic spectra, particularly for large PAHs, which are considered key sources of AIBs, and for comprehensively studying the spectroscopic properties of PAHs with extensive structural diversity.
The AIB features are believed to originate from the IR emission of PAHs excited by UV photons. Upon absorbing such a photon, a PAH becomes highly energized, with its effective temperature potentially rising to several hundred or even over a thousand kelvin. Consequently, the low-temperature spectra of PAHs mentioned earlier have limited astronomical relevance in this context. Traditionally, the Wang and Landau method has been used to incorporate temperature effects into a DFT-computed IR spectrum (Wang & Landau, 2001; Basire et al., 2009). However, as discussed above, such DFT calculations are computationally expensive. Moreover, the Wang-Landau algorithm often requires extensive sampling to ensure convergence, making it suitable mainly for studying individual, very small PAHs (Chen, 2018; Chen et al., 2018; Mackie et al., 2018b, 2021; Chakraborty et al., 2021). In contrast, the MLMD approach employed in this work explicitly incorporates temperature effects, providing a more efficient solution due to its high computational efficiency and accuracy.
To demonstrate the precision, we compared theoretical anharmonic spectra with experimental IR spectra of two PAHs, pyrene and benzo[$k$]fluoranthene, measured at different temperatures. The higher temperature gas-phase experimental spectra have been taken from the NIST Chemistry WebBook database (Linstrom & Mallard, 2001), while the lower temperature spectra come from matrix isolation experiments of Hudgins & Sandford (1998a); Hudgins & Allamandola (1999). Figure 6 shows the NIST spectra (gray curves), referred to as $T_{\text{high}}$, and the low-temperature spectra (black curves), referred to as $T_{\text{low}}$. Comparison of the two sets of experimental spectra measured at different temperatures reveals that thermal effects lead to notable changes in spectral characteristics, such as band broadening and peak shifts (indicated by dashed lines), which aligns with previous observations of Chen (2018).
[…]
The two pairs of experimental spectra shown in the top panels of Figure 6 a and b are compared with three theoretical spectra in the bottom panels, calculated using MLMD at 300 (red) and 50 K (blue), as well as by the anharmonic DFT method (green). Compared with the 50 K MLMD results, the temperature-induced band broadening and peak shift behaviors are better captured by the MLMD results at 300 K, with a low RMSE value of 0.060 for the two molecules. In contrast, the agreement with the room temperature experimental spectra is poorer for both the 50 K MLMD and anharmonic DFT results, with higher RMSE values of 0.098 and 0.072 for pyrene, and 0.069 and 0.062 for benzo[$k$]fluoranthene, respectively.
Despite the growing need to incorporate anharmonicity and temperature effects into astronomical spectroscopic analyses, driven by high-resolution infrared observations in the JWST era (Peeters et al., 2021), only a limited number of theoretical anharmonic spectra for PAHs, likely fewer than 100, are available in the literature. Typically, these studies focus on just 2-5 PAH species, most of which contain fewer than 20 carbon atoms (Esposito et al., 2024c, d; Mackie et al., 2022, 2021). This narrow scope makes it difficult to draw general conclusions, especially given the vast diversity of PAH chemical structures. Such limitations hinder the accurate interpretation of AIBs and are mainly due to the high computational cost of traditional QCCs. To address this gap, we have computed the anharmonic spectra at 50, 300, and 600 K for 1 704 molecules from version 3.20 of PAHdb, which includes neutral PAHs as large as 216 carbon atoms. This dataset is open to the public as described in the Section of Data Availability.
This dataset is expected to be a valuable resource for the spectral decomposition of AIBs and for data-driven studies investigating the precise structure-spectral relationships of PAHs. In addition, it will aid in the exploration of temperature effects. For instance, Figure 7 displays the anharmonic IR spectra of a simple mixture of the 1 704 neutral PAHs at 50 K (blue), 300 K (green), and 600 K (red). These spectra represent the sum of individual PAH spectra and are compared to the ground-state harmonic spectra from PAHdb (gray). Notably, the IR signals are significantly enhanced at high temperatures, exhibiting stronger peaks and an elevated plateau. A redshift in peak positions is also observed at high temperatures as a general trend. Furthermore, some emission bands that are weak at lower temperatures become significantly stronger at 300 and 600 K. For example, a distinct band emerges in the 3.5-6 $\upmu$m region, absent in most harmonic spectra but observed in AIBs (Jourdain de Muizon et al., 1986; Sloan et al., 1997). This feature may originate from the redshift of the C-H stretching band in superhydrogenated PAHs due to anharmonic effects (Mackie et al., 2018a; Yang et al., 2020; Sandford et al., 2013), with additional redshift and intensity enhancement likely driven by temperature effects.
Figure 7: IR spectra of a mixture of 1 704 neutral PAHs computed using MLMD at 50 K (blue), 300 K (green), and 600 K (red). For comparison, the mixture of scaled harmonic spectra from PAHdb, are shown in gray. Each spectrum is normalized to its respective maximum intensity.
IV Conclusions
We have demonstrated that, when applied to the prediction of the IR spectra of PAHs, MLMD provides predictive accuracy comparable to that of traditional quantum chemical methods but at a fraction of the computational cost. With a scaling law of approximately $N_{\text{C}}^{0.7}$, MLMD is particularly suitable for large PAHs, which are key contributors to AIBs. In contrast, traditional quantum chemical anharmonic methods scale much more steeply, approximately $N_{\text{C}}^{4.1}$. Furthermore, we have shown that MLMD is an ideal method for studying temperature effects in molecular IR spectra, which might be critical for understanding astronomical observations. By calculating the anharmonic spectra of 1 704 neutral PAHs from PAHdb at 50, 300, and 600 K, with sizes up to 216 carbon atoms, we highlight that the efficiency of MLMD enables the creation of extensive molecular anharmonic spectral datasets. These datasets could be of instrumental importance for the advancement of AI-assisted astronomical analyses of observational IR spectra in the future.
Despite the impressive efficiency of MLMD, several limitations persist. First, the current model was trained exclusively on neutral PAHs with natural chemical elements and does not account for charged molecules or isotopologues, both of which may play a significant role in astronomical contexts. This limitation stems from the scarcity of data for the IR spectra, as well as challenges in incorporating charge state and isotope information into molecular descriptors. Second, because of the data-driven nature of the approach, the model may exhibit increased uncertainty when predicting spectra for molecules that deviate substantially from those in the training dataset. We are actively working to address these challenges, with the goal of further enhancing MLMD’s capabilities for IR spectrum computation.
Data Availability
The source code and the MLMD model to calculate the anharmonic spectra of PAHs using MLMD are freely available as the supplementary information of this article. The model is trained and ready for use without the need for a further ML procedure. It will be continuously updated to improve predictive performance. The supplementary file also includes the spectral data for the 1 704 theoretically-calculated and 49 experimentally-tested PAHs.
Appendix I
The structures and chemical formulas of the 49 PAHs, whose experimentally measured IR spectra are compared with the theoretical spectra in this study, are presented in Figure 8. The spectra and the corresponding relaxed structures are obtained from version 3.2 of PAHdb (Mattioda et al., 2020).
Figure 8: Structures and chemical formulas of the 49 experimentally measured PAHs.